# Understanding Action Effects through Instrumental Empowerment in Multi-Agent Reinforcement Learning

### v1

@@ 4.2 Action Attribution
$\Phi_{i}\left(\nu\right)\coloneqq\frac{1}{T}\sum_{t=0}^{T-1}\sum_{\sigma\in\Sigma_{+}^{i}}w\cdot\Delta\nu\left(\mathcal{C}_{\sigma}^{i},s_{t}\right).$
In practice, it is computationally infeasible to consider all coalitions for large $n$ [40]. As in previous works [7], we use Monte Carlo sampling to get an unbiased estimator $\hat{\Phi}$ by sampling a new order $\sigma_{t}$ for each step $t$ and by assuming $T\gg n$, as follows:
$\hat{\Phi}_{i}\left(\nu\right)\coloneqq\frac{1}{T}\sum_{t=0}^{T-1}\cdot\Delta\nu\left(\mathcal{C}_{\sigma_{t}}^{i},s_{t}\right),\quad\sigma_{t}\sim\mathcal{U}_{\Sigma_{+}^{i}}.$ (4)
@@
Figure 3: Intermediate states $S_{t,(\cdot)}^{\sigma(\cdot)}$ matching the color of the processed action $a_{t}^{\sigma(k)}$ are used for computing the marginal contribution $\Delta\nu$. Notice that $S_{t,(1)}^{\sigma(3)}$ is used for the effects of both $a_{t}^{\sigma(1)}$ and $a_{t}^{\sigma(2)}$. Arcs entering $a_{t}^{\sigma(k)}$ are explicitly ignored due to intervention.
@@ 4.3 Characteristic Functions
which reduces to $\nu_{v}(\mathcal{C}_{\sigma}^{i},s_{t,(k)})=V(s_{t,(k)})$ in the case of centralized critics, with $V^{j}(\cdot)$ being defined in Eq. (1).
@@
As noted in the introduction, access to many future states is instrumentally valuable. The size of future accessible states can be quantified by $\lvert\mathcal{S}_{+}\rvert$, where $\mathcal{S}_{+}=\{s_{t+1}\in\mathcal{S}:P(s_{t+1}\mid s_{t})>0\}$. Based on this, Franzmeyer et al. [10] empirically validated the use of $H(A_{t}\mid s_{t})$ as a proxy for $\lvert\mathcal{S}_{+}\rvert$ to capture an agent’s future choices. Similarly, Watson et al. [44] introduced information-theoretic SVs via $v=H(Y\mid z_{\mathcal{K}})$ to measure predictive uncertainty in ML, with $Y$ being the output. Building on both, we argue that in some multi-agent settings, it may be more valuable for an agent to be certain about what to do, i.e., committed to a specific strategy. Since $H(A_{t,(k+1)}^{i}\mid s_{t,(k)})\leq\log|\mathcal{A}^{i}|$ [13] quantifies the spread of $\pi^{i}(\cdot\mid s_{t,(k)})$, we define its complement to this bound as a measure of peakedness: $\mathrm{H}(A_{t,(k+1)}^{j})\coloneqq\log\lvert\mathcal{A}^{j}\rvert-H(A_{t,(k+1)}^{j}\mid s_{t,(k)})$.
@@
The characteristic peak-based function is defined as:
$\nu_{p}(\mathcal{C}_{\sigma}^{i},s_{t,(k)})\coloneqq\sum_{j\in\mathcal{C}_{\sigma}^{i}}\mathrm{H}\left(A_{t,(k+1)}^{j}\right)$
@@
$\mathrm{JSD}(p\mid\mid q)=\frac{1}{2}D_{KL}(p\mid\mid u)+\frac{1}{2}D_{KL}(q\mid\mid u).$ (7)
@@
Using base-2 $\log$, $\mathrm{JSD}(p\|q)$ is bounded in $[0,1]$, where 0 indicates identical distributions and $1$ maximal divergence. To instead express similarity, we define $J(p\|q)\coloneqq 1-\mathrm{JSD}(p\|q)$.
@@
The characteristic consensus-based function is:
$\nu_{c}(\mathcal{C}_{\sigma}^{i},s_{t,(k)})\coloneqq$ $\sum_{j\in\mathcal{C}_{\sigma}^{i}}\bigg{[}\overbrace{J\left(\pi^{i}\left(\cdot\mid s^{(j)}_{t,(k)}\right)\mid\mid\pi^{j}\left(\cdot\mid s^{(j)}_{t,(k)}\right)\right)}^{other}$
$+\underbrace{J\left(\pi^{i}\left(\cdot\mid s^{(i)}_{t,(k)}\right)\mid\mid\pi^{j}\left(\cdot\mid s^{(i)}_{t,(k)}\right)\right)}_{self}\bigg{]}.$
If only one of the consensus terms is used, we denote others-consensus by $\nu_{co}$ and self-consensus by $\nu_{cs}$. For dissimilarity, we use $\mathrm{JSD}$ in place of $J$ in $\nu_{c}$ and term it $\nu_{d}$, $\nu_{do}$ and $\nu_{ds}$ respectively.
@@ 4.4 Instrumental Empowerment
The advantage of agent $j$ is defined by $\boldsymbol{\Lambda}^{j}(s_{t},\boldsymbol{a}_{t})\coloneqq Q^{j}(s_{t},\boldsymbol{a}_{t})-V^{j}(s_{t})=\mathbb{E}_{s_{t+1}\sim P}[R^{j}+\gamma V^{j}(s_{t+1})]-V^{j}(s_{t})$.
@@
For each intermediate state $s_{t,(k)}$, $\Delta\nu_{v}$ measures the sampled advantage for players in $\mathcal{C}_{\sigma}^{i}$ attributable to action $a_{t}^{i}$:
$\Delta\nu_{v}=\sum_{j\in\mathcal{C}_{\sigma}^{i}}V^{j}(s_{t,(k)})-V^{j}(s_{t,(k-1)})=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\Lambda^{j}(s_{t,(k)},a_{t}^{i})$
@@
The empowerment function [16] quantifies an agent’s maximal potential causal influence on future states, defined as $\mathcal{E}(S_{t})=\max_{p(a_{t}\mid s_{t})}\left[H(S_{t+1}\mid S_{t})-H(S_{t+1}\mid S_{t},A_{t})\right]$.
@@
For each intermediate state $s_{t,(k)}$, $\Delta\nu_{p}$ quantifies the instrumental empowerment as the increase in decision certainty of players in $\mathcal{C}_{\sigma}^{i}$ attributable to action $a_{t}^{i}$:
$\Delta\nu_{p}=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\mathrm{H}\left(A_{t,(k+1)}^{j}\right)-\mathrm{H}\left(A_{t,(k)}^{j}\right)=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\mathcal{E}^{j}(s_{t,(k-1)})$
@@ 5 Experiments
To assess whether the intuition from Fig. 1 and our ICV method generalize to more complex settings, we conduct empirical validation in three environments. [^1] Python code available at github.com/icv-marl. Each illustrates distinct applications of ICVs online and offline for extracting behavioral insights and identifying credit assignment in the absence of rewards. We examine whether $\Phi(\nu_{p})$ and $\Phi(\nu_{c})$ can serve as proxies for $\Phi(\nu_{v})$. As the validity of using the value-based function $\nu_{v}$ in $\Phi(\nu_{v})$ as a feature contribution predictor has been proven in [3], it serves as our baseline. We compute ICVs over $M$ episodes and report relative ICV values by $i)$ normalizing all characteristic functions $\nu$, and ii)ii) dividing each $\Phi(\cdot)$ by $\kappa=\max\Phi(\cdot)$ to scale results close to $1$, as some contributions may cancel out and yield low contributions on average. We assume convergence of the learning process, i.e., no further performance improvement, but note that our method is applicable at any point during training.
@@ 5.1 Cooperative Game
For the fully cooperative case, we consider the environment Level-based Foraging (LBF), illustrated in Fig. 4(a). Agents and apples are assigned levels, and picking up an apple requires the sum of the levels of nearby agents to match the apple’s level, followed by a simultaneous load action. Each agent uses an individual centralized critic, and our ICV method is applied online. We first validate the correspondence between value $V^{i}$ and peak $\mathrm{H}^{i}$ increase in Fig. 4. In the depicted initial state (left), agent 1 moves to the right (green arrow), thereby increasing both its own and its teammates’ certainty in action selection, while also increasing the value for all, which is shown in the right plot. This effect arises because agent 1 positions itself below the apple, therefore clarifying to others where they should move to.
@@ 5.2 Mixed-Motive Game
For the mixed-motive setting, we use the Multi-Particle Environment (MPE) Tag, shown in Fig. 7(a), where predators (red) aim to catch a prey (green). Agents have individual centralized critics, and ICV computation is performed offline, although online computation would be feasible since each agent is processed independently. In Fig. 7(b), we analyze agent 1 (prey) and again verify the correspondence between its value $V^{1}$ and action determinism $\mathrm{H}^{1}$ over a single episode. The results indicate a reasonable alignment between value and action certainty. We further assess agent 1’s dissimilarity to its opponents (predators) by setting $\mathcal{C}_{\sigma}^{i}$ accordingly and measuring $\nu_{d}$, denoted as $\mathrm{J}^{1}$. Agent 1 shows high dissimilarity throughout the episode which is consistent with the intuition that a prey should act contrary to predators, ideally avoiding movement toward them.
(a) MPE environment.
$V^{1}$   $\mathrm{H}^{1}$   $\mathrm{J}^{1}$
@@
Figure 7: Comparing normalized value $V^{1}$, $\mathrm{H}^{1}$ and $\mathrm{J}^{1}$ on the right plot along y-axis over time steps $t$ along x-axis of an MPE prey.
We measured the ICVs over $M=500$ episodes and normalized them using $\kappa=0.08$. Due to the symmetry of the game, we focus on the contributions of a single adversary (predator, denoted as Adv) and a single agent (prey, denoted as Ag). For both, we partition their coalitions and evaluate the effects on teammates (-t) and opponents (-o), where $\nu_{c}$ is used for teammates, while dissimilarity $\nu_{d}$ is applied to opponents. For predators, this means: in Adv-t, $\nu_{c}$ captures similarity among predators and $\nu_{d}$ measures dissimilarity between predators and prey. In Adv-o, $\nu_{c}$ reflects similarity among prey, and $\nu_{d}$ their dissimilarity with predators. The same interpretation applies to the contributions of prey (Ag).
Figure 8 presents the results, where proximity to others is the main effect. The prey’s contributions align with our expectations: average value effects $\hat{\Phi}(\nu_{v})$ correspond to action determinism $\hat{\Phi}(\nu_{p})$ and alignment with co-player’s preferences (Ag-t, $\hat{\Phi}(\nu_{c})$), but interestingly, they contribute to less diversity toward predators $\hat{\Phi}(\nu_{d})$. In contrast, predators benefit from reducing action determinism and keeping their choices open when engaging prey, as indicated by increased $\hat{\Phi}(\nu_{c})$ but decreased $\hat{\Phi}(\nu_{p})$. They also tend to behave slightly more similar to each other (Adv-t, $\hat{\Phi}(\nu_{c})$) while promoting diversity against opponents ($\hat{\Phi}(\nu_{d})$). In Adv-o, high $\hat{\Phi}(\nu_{p})$ and low $\hat{\Phi}(\nu_{v})$ indicate prey become more deterministic (e.g., fleeing) as predators approach. Their effect also renders prey reacting differently (Adv-o, $\hat{\Phi}(\nu_{c})$). Overall, compared to Fig. 7(b), the ICVs suggest that $\hat{\Phi}(\nu_{p})$ can signal both positive and negative contributions.

### v2

@@ 4.2 Action Attribution
$\Phi_{i}\left(\nu\right)\coloneqq\frac{1}{T}\sum_{t=0}^{T-1}\sum_{\sigma\in\Sigma_{+}^{i}}w\cdot\Delta\nu\left(\mathcal{C}_{\sigma}^{i},s_{t,(k)}\right).$
For standard SV computation, the time complexity is $O(2^{n})$. Our method increases this by a factor of $n$, since each ordering requires $n$ computations of marginal contributions. Explainability methods like ours are usually applied in simulation to gain insights prior to real-world deployment, where higher computational costs are acceptable. Nevertheless, in practice it remains computationally infeasible to consider all coalitions for large $n$ [36]. As in previous works [6], we use Monte Carlo sampling for orders $\sigma_{t}$ to get an unbiased estimator $\hat{\Phi}$ for each step $t$ and by assuming $T\gg n$, as follows:
$\hat{\Phi}_{i}\left(\nu\right)\coloneqq\frac{1}{T}\sum_{t=0}^{T-1}\cdot\Delta\nu\left(\mathcal{C}_{\sigma_{t}}^{i},s_{t,(k)}\right),\quad\sigma_{t}\sim\mathcal{U}_{\Sigma_{+}^{i}}.$ (4)
@@
Figure 3: Intermediate states $S_{t,(\cdot)}^{\sigma(\cdot)}$ used for computing the marginal contributions $\Delta\nu$ match the color of the processed action $a_{t}^{\sigma(k)}$. Notice that $S_{t,(1)}^{\sigma(3)}$ is used for the effects of both $a_{t}^{\sigma(1)}$ and $a_{t}^{\sigma(2)}$. Arcs entering $a_{t}^{\sigma(k)}$ are explicitly ignored due to $\mathrm{do}$-intervention.
@@ 4.3 Characteristic Functions
with $V^{j}(\cdot)$ being defined in Eq. (1). This reduces to $\nu_{v}(\mathcal{C}_{\sigma}^{i},s_{t,(k)})=V(s_{t,(k)})$ in the case of centralized critics.
@@
As noted in Sec.1, access to many future states is instrumentally valuable. The size of future accessible states can be quantified by $\lvert\mathcal{S}_{+}\rvert$, where $\mathcal{S}_{+}=\{s_{t+1}\in\mathcal{S}:P(s_{t+1}\mid s_{t})>0\}$. Based on this, Franzmeyer et al. [9] empirically validated the use of $H(A_{t}\mid s_{t})$ as a proxy for $\lvert\mathcal{S}_{+}\rvert$ to capture an agent’s future choices. Similarly, Watson et al. [40] introduced information-theoretic SVs via $v=H(Y\mid z_{\mathcal{K}})$ to measure predictive uncertainty in ML, with $Y$ being the output. Building on both, we argue that in some multi-agent settings, it may be more valuable for an agent to be certain about what to do, i.e., committed to a specific strategy.
@@
Since $H(A_{t,(k+1)}^{i}\mid s_{t,(k)})\leq\log|\mathcal{A}^{i}|$ [12] quantifies the spread of $\pi^{i}(\cdot\mid s_{t,(k)})$, we define the complement to this bound as a measure of decision certainty, denoted as peakedness:
$\mathcal{H}(A_{t,(k+1)}^{j})\coloneqq\log\lvert\mathcal{A}^{j}\rvert-H(A_{t,(k+1)}^{j}\mid s_{t,(k)}).$
The characteristic peak-based function is then given as:
$\nu_{p}(\mathcal{C}_{\sigma}^{i},s_{t,(k)})\coloneqq\sum_{j\in\mathcal{C}_{\sigma}^{i}}\mathcal{H}\left(A_{t,(k+1)}^{j}\right).$
@@
$\bar{\mathcal{J}}(p\mid\mid q)\coloneqq\mathrm{JSD}(p\mid\mid q)=\frac{1}{2}D_{KL}(p\mid\mid u)+\frac{1}{2}D_{KL}(q\mid\mid u).$ (7)
@@
Using base-2 $\log$, $\bar{\mathcal{J}}(p\|q)$ is bounded in $[0,1]$, where 0 indicates identical distributions and $1$ maximal divergence. To instead express similarity, we define $\mathcal{J}(p\|q)\coloneqq 1-\bar{\mathcal{J}}(p\|q)$. The characteristic consensus-based function is thus given by:
$\nu_{c}(\mathcal{C}_{\sigma}^{i},s_{t,(k)})\coloneqq$ $\sum_{j\in\mathcal{C}_{\sigma}^{i}}\bigg{[}\overbrace{\mathcal{J}\left(\pi^{i}\left(\cdot\mid s^{(j)}_{t,(k)}\right)\mid\mid\pi^{j}\left(\cdot\mid s^{(j)}_{t,(k)}\right)\right)}^{other}$
$+\underbrace{\mathcal{J}\left(\pi^{i}\left(\cdot\mid s^{(i)}_{t,(k)}\right)\mid\mid\pi^{j}\left(\cdot\mid s^{(i)}_{t,(k)}\right)\right)}_{self}\bigg{]}.$
If only one of the consensus terms is used, we denote others-consensus by $\nu_{co}$ and self-consensus by $\nu_{cs}$. For dissimilarity, we use $\bar{\mathcal{J}}$ in place of $\mathcal{J}$ in $\nu_{c}$ and term it $\nu_{d}$, $\nu_{do}$ and $\nu_{ds}$ respectively.
@@ 4.4 Instrumental Empowerment
The advantage of agent $j$ is given by $\boldsymbol{\Lambda}^{j}(s_{t},\boldsymbol{a}_{t})\coloneqq\mathbb{E}_{s_{t+1}\sim P(\cdot|s_{t},\boldsymbol{a}_{t})}[R^{j}(s_{t},\boldsymbol{a}_{t},s_{t+1})+\gamma V^{j}(s_{t+1})]-V^{j}(s_{t})$. We define the sampled advantage as:
$\boldsymbol{\hat{\Lambda}}^{j}(s_{t},\boldsymbol{a}_{t},s_{t+1})\coloneqq R^{j}(s_{t},\boldsymbol{a}_{t},s_{t+1})+\gamma V^{j}(s_{t+1})-V^{j}(s_{t}).$
@@
Given $V^{j}(s)$ as defined by Eq. (1) that is learned under $\mathcal{G}_{m}$ dynamics, and employing it as a static predictor within $\mathcal{G}_{s}$, then for each transition of intermediate states $s_{t,(k-1)}\rightarrow s_{t,(k)}$, the value-based marginal contribution $\Delta\nu_{v}$ measures the undiscounted sampled advantage for players $j\in\mathcal{C}_{\sigma}^{i}$ attributable to agent $i$:
$\Delta\nu_{v}=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\hat{\Lambda}^{j}(s_{t,(k-1)},a_{t}^{i},s_{t,(k)}).$
@@
The ordinary empowerment function [14] quantifies an agent’s maximal potential causal influence on future states. We define the instrumental empowerment of the agent $i=\sigma(k)$ as the potential influence on future actions of agents $j\in\mathcal{C}^{i}_{\sigma}$ as follows:
$\mathcal{E}^{j}(s_{t,(k-1)},i)\coloneqq\max_{\pi^{i}\in(\mathcal{P}(\mathcal{A}^{i}))^{\mathcal{S}}}\,\mathrm{I}(A_{t,(k)}^{j};a_{t}^{i}\mid s_{t,(k-1)}).$
@@
For each intermediate state $s_{t,(k)}\in\mathcal{S}$, the entropy-based marginal contribution $\Delta\nu_{p}$ measures the mutual information $\mathrm{I}(A_{t,(k)}^{j};a_{t}^{i}\mid s_{t,(k-1)})$ as the certainty increase in decision-making of players $j\in\mathcal{C}_{\sigma}^{i}$ attributable to agent $i$’s action $a_{t}^{i}$:
$\Delta\nu_{p}=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\mathrm{I}(A_{t,(k)}^{j};a_{t}^{i}\mid s_{t,(k-1)}).$
If moreover the acting policy $\pi^{i}$ maximizes each conditional mutual information, then $\Delta\nu_{p}$ induces the instrumental empowerment:
$\Delta\nu_{p}=\sum_{j\in\mathcal{C}_{\sigma}^{i}}\mathcal{E}^{j}(s_{t,(k-1)},i).$
@@ 5 Experiments
To assess whether the intuition from Fig. 1 and our ICV method generalize to more complex settings, we conduct empirical validation in three environments. [^1] Python code available at github.com/icv-marl. Each illustrates distinct applications of ICVs online and offline for extracting behavioral insights and identifying credit assignment in the absence of rewards. We examine whether $\Phi(\nu_{p})$ and $\Phi(\nu_{c})$ can serve as proxies for $\Phi(\nu_{v})$. Thus, our reward-free method can be seen as an extension of the reward-based approaches in [3, 24, 32], particularly since the validity of using the value-based function $\Phi(\nu_{v})$ as a feature contribution predictor has been proven in [3]. We compute ICVs over $M$ episodes and report relative ICV values by $i)$ normalizing all characteristic functions $\nu$, and ii)ii) dividing each $\Phi_{i}(\cdot)$ by $\kappa=\max_{i,j}\Phi_{i}(\nu_{j})$ to scale results close to $1$, as some contributions may cancel out and yield low contributions on average. We consider pre-trained models, with training similar to [26], and assume convergence of the learning process but note that our method is applicable at any point during training.
@@ 5.1 Cooperative Game
We consider the fully cooperative task in Level-based Foraging (LBF), illustrated in Fig. 4(a), where the numbering of agents and apples serve as instance identifiers (rather than level indicators). All agents must surround the same apple on separate cells and then perform a simultaneous load action to receive a reward.
Each agent uses an individual centralized critic, and our ICV method is applied online. We first validate the correspondence between value $V^{i}$ and peak $\mathcal{H}^{i}$ increase in Fig. 4. In the depicted initial state (left), agent 1 moves to the right (green arrow), thereby increasing both its own and its teammates’ certainty in action selection, while also increasing the value for all, which is shown in the right plot. This effect arises because agent 1 positions itself below the apple, therefore clarifying to others where they should move to.
@@ 5.2 Mixed-Motive Game
Here, we use the mixed-motive Tag game (Fig. 7(a)) of Multi-Particle Environment (MPE), where predators (red) aim to catch a prey (green). Agents have individual critics, and ICVs are computed offline. In Fig. 7(b), we analyze agent 1 (prey) and again verify the correspondence between its value $V^{1}$ and action determinism $\mathcal{H}^{1}$ over a single episode. The results indicate a reasonable alignment between value and action certainty. We further assess agent 1’s dissimilarity to its opponents (predators) by setting $\mathcal{C}_{\sigma}^{i}$ accordingly and measuring $\nu_{d}$, denoted as $\bar{\mathcal{J}}^{1}$. Agent 1 shows high dissimilarity throughout the episode which is consistent with the intuition that a prey should act contrary to predators, ideally avoiding movement toward them.
(a) MPE tag.
$V^{1}$   $\mathcal{H}^{1}$   $\bar{\mathcal{J}}^{1}$
@@
Figure 7: Comparing normalized value $V^{1}$, $\mathcal{H}^{1}$ and $\bar{\mathcal{J}}^{1}$ on the right plot along y-axis over time steps $t$ along x-axis of an MPE prey.
We measured the ICVs over $M=500$ episodes and normalized them using $\kappa=0.08$. Due to game symmetry, we focus on a single adversary (predator, denoted as adv) and a single agent (prey, denoted as ag). For both, we partition their coalitions and evaluate the effects on teammates (-t) and opponents (-o), where $\nu_{c}$ is used for teammates, while dissimilarity $\nu_{d}$ is applied to opponents. For predators, this means: in adv-t, $\nu_{c}$ captures similarity among predators and $\nu_{d}$ measures dissimilarity between predators and prey. In adv-o, $\nu_{c}$ reflects similarity among prey, and $\nu_{d}$ their dissimilarity with predators. The same interpretation applies to prey (ag).
Figure 8 presents the results, where proximity to others is the main effect. The prey’s contributions align with our expectations: average value effects $\hat{\Phi}(\nu_{v})$ correspond to action determinism $\hat{\Phi}(\nu_{p})$ and alignment with co-player’s preferences (ag-t, $\hat{\Phi}(\nu_{c})$), but interestingly, they contribute to less diversity toward predators $\hat{\Phi}(\nu_{d})$. In contrast, predators benefit from reducing action determinism and keeping their choices open when engaging prey, as indicated by increased $\hat{\Phi}(\nu_{c})$ but decreased $\hat{\Phi}(\nu_{p})$. They tend to behave slightly different (adv-t, $\hat{\Phi}(\nu_{c})$) while also promoting diversity against opponents ($\hat{\Phi}(\nu_{d})$). In adv-o, high $\hat{\Phi}(\nu_{p})$ and low $\hat{\Phi}(\nu_{v})$ indicate prey become more deterministic (e.g., fleeing) as predators approach. Their effect also renders prey reacting differently (adv-o, $\hat{\Phi}(\nu_{c})$). Overall, compared to Fig. 7(b), the ICVs suggest that $\hat{\Phi}(\nu_{p})$ largely behaves as expected but may signal both positive and negative contributions.